# Discretized Quantum Exhaustive Search for Variational Quantum Algorithms

### v1

@@ Abstract
Quantum computers promise a great computational advantage over classical computers, yet currently available quantum devices have only a limited amount of qubits and a high level of noise, limiting the size of problems that can be solved accurately with those devices. Variational Quantum Algorithms (VQAs) have emerged as a leading strategy to address these limitations by optimizing cost functions based on measurement results of shallow-depth circuits. However, the optimization process usually suffers from severe trainability issues as a result of the exponentially large search space, mainly local minima and barren plateaus.
Here we propose a novel method that can improve variational quantum algorithms — “discretized quantum exhaustive search”. On classical computers, exhaustive search, also named brute force, solves small-size NP complete and NP hard problems. Exhaustive search and efficient partial exhaustive search help designing heuristics and exact algorithms for solving larger-size problems by finding easy subcases or good approximations. We adopt this method to the quantum domain, by relying on mutually unbiased bases for the $2^{n}$-dimensional Hilbert space. We define a discretized quantum exhaustive search that works well for small size problems. We provide an example of an efficient partial discretized quantum exhaustive search for larger-size problems, in order to extend classical tools to the quantum computing domain, for near future and far future goals. Our method enables obtaining intuition on NP-complete and NP-hard problems as well as on Quantum Merlin Arthur (QMA)-complete and QMA-hard problems.
@@ 1 Introduction
To make use of these NISQ machines, a class of hybrid quantum-classical algorithms that seek to leverage the relative strengths of quantum and classical computers is being developed. The most common example of such algorithms is Variational Quantum Algorithms (VQAs)[12]. VQAs use a quantum computer to prepare a short-depth Parameterized Quantum Circuit (PQC) representing a trial solution or ansatz to the problem at hand. Measurements of a final quantum state are used to calculate a cost function, which is then minimized on a classical computer to estimate the problem solution. Prominent examples of VQAs include the Variational Quantum Eigensolver (VQE) for quantum chemistry and materials applications [35, 24], Quantum Approximate Optimization Algorithm (QAOA) for combinatorial optimization problems [17], and Variational Quantum Linear Solver (VQLS) for linear algebra problems [9]. More details about these methods are provided in the appendix.
One of the main challenges associated with effective VQA implementations is related to the design of a suitable PQC/ansatz that balances expressibility and noise, while avoiding exponentially vanishing gradients of the cost function, referred to as the barren plateau (BP) problem [29].
Designing better ansatzes for VQAs is an active research area that has drawn much attention lately. There have been a number of attempts [44] to improve different ansatz approaches, including ADAPT-VQE [19] and Noise-Adaptive Search (QuantumNAS) [46].
Another important research area is developing methods for finding good initial state candidates for an ansatz. Finding a good initial guess for the optimizer is very important, as avoiding BP using smart ansatz engineering without drastically changing the problem’s search space might be impossible [11]. A good initial guess can make the difference between converging to the optimal solution or leading the optimizer to a local minimum trap. A good initial guess becomes more important as the problem size increases and the problem’s Hilbert space and the search spaces become exponentially larger. This increase in search size usually creates exponentially small gradients in the cost function landscape, making the optimization process almost impossible.
@@
We suggest to use a set of discrete points that are equally spread all over the search space of the problem. We define “discretized quantum exhaustive search” (DQES) in analogy to the use of standard exhaustive search in modern computer science.
@@
In addition, no method to sample the entire Hilbert space using evenly spaced states was defined. Using evenly spread states is important for gaining knowledge in a well-organized way about the landscape of solutions to a problem, and for drawing energy curves once applying optimizer for VQA, especially with the goal of bypassing barren plateaus.
@@
The Clifford states have many benefits, but they are not suitable for every cost function, as the Clifford gates are stabilized by the Pauli operators. The expectation value of a Clifford state against a Pauli operator can lead to either an eigenvalue of $+1$ or $-1$, in case this Clifford is indeed stabilized by that particular Pauli operator, or to 00. For each Pauli operator, there are exponentially more Clifford states whose expectation value is 00 compared to those whose expectation value is either $+1$ or $-1$. Many common VQAs use Pauli expectation values to evaluate their cost function, as it can be implemented very efficiently and easily on a quantum computer.
The main goal of this research is to define novel methods to improve VQA ansatzes and initial states, leading to better accuracy and faster convergence rates. This goal is achieved by a few different novel methods -
  • Describing a new method to sample and explore VQAs’ problem space, using the defined DQES. This method allows one to understand the problem’s Hilbert space better, creating better guesses for initial parameter choices, and designing methods to avoid BP and local minima. This method is described in Sections 2.1 and 2.2. Examples of using this method are presented in Section 3.
  • Describing an alternative heuristic method to explore the VQA’s problem space, using an efficient approach. This method aims to bring a heuristical equivalent of the above method for larger problem instances while remaining computationally feasible. This method is presented in Section 2.4. Examples of using this method are also presented in Section 3.
1.1 Mutually Unbiased Bases
Mutually unbiased bases (MUBs) are orthonormal bases in a complex vector space that provide complementary and non-redundant information about the quantum states of different bases if measured [41, 47, 5, 16]. When two or more bases are mutually unbiased, it means that the measurement outcomes in one basis provide no information about the outcomes in another basis, and if you have a quantum state prepared in one basis and you measure it in another basis, the measurement outcomes are uncorrelated.
@@ 1.1 Mutually Unbiased Bases
2 Methods
VQAs search for the optimal solution of the given problem using the search space imposed by the chosen ansatz. The Hilbert space of the problem is determined only by the problem definition and encoding and is not subject to changes by replacing the ansatz. The dimension of this Hilbert space is exponential in the number of qubits required to encode the problem into a quantum computer. Finding the optimal solution, or even a good approximation to it in such a huge space is usually a very hard task. In many search spaces, the gradients of the cost function become exponentially small in most areas of the search space, a phenomenon called barren plateaus (BP), and most of the states that have significant gradients are concentrated in a small area [29, 3]. The concentration of most of the cost function values that have a significant difference between them in a small area makes the search hard also for gradient-free optimizers [2].
@@ 2 Methods
In our research, we propose methods to learn and gain insights about the values and structure of the cost function in the problem’s Hilbert space, in a way that is agnostic to the ansatz that will be used to try to solve the problem. Using such methods can help researchers design ansatzes, optimization algorithms, and initial guesses that can avoid BPs.
@@
2.1 Discretized Quantum Exhaustive Search
@@
Here we define DQES which is useful for small size problems, and efficient partial DQES (for large-size problems), in order to extend these classical tools into the quantum computing domain, for near future and far future goals. We do so by relying on mutually unbiased bases for $2^{n}$-size Hilbert space. For non-partial DQES, we limit ourselves to small $n$.
Our method enables obtaining intuition on NP-complete and NP-hard problems as well as on Quantum Merlin Arthur (QMA)-complete and QMA-hard problems.
@@
NP problems can be solved in polynomial time by a classical computer once a relevant and correct hint is given from an external source. In a similar way, the QMA complexity class contains problems that can be solved by a quantum computer in polynomial time once a relevant hint is given by an external source. We notice that the method of VQE resembles the idea of using a hint, yet instead of obtaining the correct hint as in QMA problems, clever scientists guess a sequencial set of hints. When a quantum computer tries to solve an NP problem of a small size, it is possible to test all possible classical hints via the exhaustive search, to gain knowledge on the structure of the problem, and design heurstics for larger size problems.
@@
As looking over all the possible states as possible hints is not possible — for a quantum space — we look for an approximated method relying on evenly spread hints, that can resemble the classical exhaustive search.
@@
Searching an exponential number of different states, which are spread evenly all over the problem’s relevant Hilbert space, for small size problems, may help designing heuristics that can then be used for larger-size problems. Specifically, it may help studying the energy landscape of problems hence the locations of barren plateaus and local minima.
2.2 MUBs for Exploring The Cost Function Landscape
@@
The MUB states of a full set of MUBs have unique properties that make them good for extracting information for purposes such as state tomography, and process tomography. In a similar manner, we suggest that these states can be used to extract information about the landscape of a cost function of a given problem, if used accordingly. A full set of MUB states can be a good candidate as states for a DQES for small-size problems. There is an exponential number of states that are spread all over the Hilbert space, and their information is, in some sense, mutually exclusive.
@@
By measuring the cost function value of a given problem at all these states, one can get evenly-spread (through space) information about the structure of the cost function. Such information can be properties of states that give similar values, hinting that there could be barren plateaus areas in the cost function landscape, properties of states that potentially present an area with local minima, and properties of states that have a larger variety of results, pointing to areas that might be interesting to search for the optimal result.
2.3 Circuit Representation of MUB States
@@
The computation space of a quantum device consisting of $n$ qubits is of dimension $2^{n}$. It has been shown [5, 47, 26, 16] that for such a space there exists a set of $12^{n}+$ 1 orthonormal bases which are mutually unbiased. Finding such a set of mutually unbiased bases for any number of qubits is not a trivial task. Few methods have been proposed to find a full set of unbiased bases, containing $12^{n}+$ 1 orthonormal bases [27, 6, 5, 47]. As far as we could figure out, the computation cost of any such known algorithm to find orthonormal bases which are part of a full MUB set, is exponential in the number of qubits. For small $n$ we noticed one research that found efficient ways to implement these MUBs in the lab using adjusted measurement equipment [21]. We need to have a quantum circuit that prepares each state in those bases. Known algorithms to find a quantum circuit that creates an arbitrary given quantum state take exponential time to compute in the worst case and might produce a very deep quantum circuit, therefore usually can not be used in the VQA context.
@@
Here we investigate two different circuit representations of MUB states, both are sufficiently shallow for our demonstrations. In the first representation, each basis in the MUB set is defined by a quantum circuit that transforms a state in the computational basis to a state in that basis. We have manually found such shallow circuits for each basis in a full set of MUBs of 2 and 3 qubits, see Fig.1. This method provides only a discrete map, yet it may be used for finding an optimal initial state (or a set of states), that can then be used as the initial state(s) of a parametrized ansatz controlled by an optimizer as in all usages of VQA.
@@
In contrast, the second representation relies directly on a chosen ansatz circuit. In this representation, each MUB state is defined by the parameters vector of an ansatz which prepares the state, hence each MUB state can then directly be used by utilizing the optimizer to modify the parameters. When using that method, the chosen ansatz must be expressive enough to allow the preparation of each of the states of the MUBs. In Fig. 1, subfigure (c), we give a hardware-efficient ansatz circuit for generating 2-qubit MUBs. When acting on the computaitonal state $|{i}\rangle$, it gives the i’th element of the MUB basis. Note that the circuit in 1 (c) can be used to generate all MUB states directly by adding a single layer of NOT gates (where needed) at the start of the circuit, if the initial state is always the all-zeros state.
@@
2.4 Efficient Partial Discretized Quantum Exhaustive Search
Another tool we suggest and demonstrate via simple examples is an efficient partial DQES. Such a search may also provide heuristics and potentially lead to exact algorithms, for solving larger-size problems, by finding easy subcases or better than classical approximations.
Relative to existing VQA methods that rely on a single initial state or just a few initial states, the method we suggest relies on a polynomial number of initial states. To obtain a scalable method we let some qubits be much more dominant than others in the cost function of a VQA problem, yet we do not assume that the dominance of each qubit is known a-priori. We suggest a method to search possible initial states by using a specific set of partial-MUB states.
Our chosen partial MUB set is defined as follows. Let $n$ denote the number of qubits in the problem. Let $K$ be a constant which we call the “partial MUB constant”. A partial-MUB state is a tensor product between $n-K$ qubits that form any single state (e.g., WLOG, the single state may be the all-zero state in the computational basis), and $K$ qubits that form a state that is part of one of the bases in a full MUB set. For small values of $K$ we may generate all MUB states. For the scalablity proof we shall not assume tha $K$ is small.
For a constant $K$, The number of different options to choose the qubits splitting is (${{n}\choose{K}}=O\left(n^{K}\right)$. The number of MUB states for $K$ qubits is $(2^{K}+1)\cdot 2^{K}=(2^{K})^{2}+2^{K}$, leaving us with a total number of
(${{n}\choose{K}}\cdot((2^{K})^{2}+2^{K})=O(n^{K}\cdot 4^{K})=O(n^{K})$ (2)

### v2

@@ Abstract
Quantum computers promise a great computational advantage over classical computers, yet currently available quantum devices have only a limited amount of qubits and a high level of noise, limiting the size of problems that can be solved accurately with those devices. Variational Quantum Algorithms (VQAs) have emerged as a leading strategy to address these limitations by optimizing cost functions based on measurement results of shallow-depth circuits. However, the optimization process usually suffers from severe convergence issues as a result of the exponentially large search space — mainly in the form of local minima and barren plateaus.
Here we propose a novel method that can improve variational quantum algorithms — “Discretized Quantum Exhaustive Search”. On classical computers, exhaustive search, also named brute force, solves small-size NP complete and NP hard problems. Exhaustive search helps design heuristics and exact algorithms for solving larger-size problems by finding easy subcases or good approximations. We adopt this method to the quantum domain, by relying on mutually unbiased bases for the $2^{n}$-dimensional Hilbert space. We define a discretized quantum exhaustive search that works well for small size problems. We provide an example of an efficient partial discretized quantum exhaustive search for larger-size problems, for near future and far future goals. Our method enables obtaining intuition on NP-complete and NP-hard problems as well as on Quantum Merlin Arthur (QMA)-complete and QMA-hard problems.
@@ I Introduction
To make use of these NISQ machines, a class of hybrid quantum-classical algorithms that seek to leverage the relative strengths of quantum and classical computers together is being developed. The most common example of such algorithms is Variational Quantum Algorithms (VQAs)[14]. VQAs use a quantum computer to prepare a short-depth Parameterized Quantum Circuit (PQC) representing a trial solution, also known as an ansatz, to the problem at hand. Measurements of a final quantum state are used to evaluate a cost function, which is then minimized on a classical computer to estimate the problem solution. Prominent examples of VQAs include the Variational Quantum Eigensolver (VQE) for quantum chemistry and materials applications [44, 26], Quantum Approximate Optimization Algorithm (QAOA) for combinatorial optimization problems [19], and Variational Quantum Linear Solver (VQLS) for linear algebra problems [11]. More details about these methods are provided in the appendices A,C,D and E.
One of the main challenges associated with effective VQA implementations is related to the design of a suitable PQC/ansatz that balances expressibility and noise, while avoiding exponentially vanishing gradients of the cost function, referred to as the barren plateaus problem [35].
Designing better ansatzes for VQAs is an active research area that has drawn much attention lately. There have been a number of attempts [53] to improve different ansatz approaches, including ADAPT-VQE [21], Differentiable Quantum Architecture Search [57], Pulse based ansatz [37, 38, 5, 31, 32] and Noise-Adaptive Search (QuantumNAS) [55].
According to [13], it might be impossible to avoid barren plateaus using smart ansatz engineering without drastically changing the problem’s search space. On the other hand, finding a good initial guess for the optimizer can possibly help avoid barren plateaus. A good initial guess can make the difference between converging to the optimal solution or leading the optimizer to a local minimum trap. A good initial guess becomes more important as the problem size increases and the problem’s Hilbert space and the search spaces become exponentially larger. This increase in search size usually creates exponentially small gradients in the cost function landscape, making the optimization process almost impossible.
@@
We suggest to use a set of discrete points that are equally spread all over the search space of the problem. In Section II.2, we define “discretized quantum exhaustive search” (DQES) in analogy to the use of standard exhaustive search in modern computer science.
@@
In addition, no method to sample the entire Hilbert space using evenly spaced states was defined. Using evenly spread states is important for gaining knowledge in a well-organized way about the landscape of solutions to a problem, and for drawing energy curves as a function of the optimziation steps, once applying an optimizer for VQA, especially with the goal of bypassing barren plateaus.
@@
The Clifford states have many benefits, but they are not suitable for every cost function, as the Clifford gates are stabilized by the Pauli operators. The expectation value of a Clifford state against a Pauli operator can lead to either an eigenvalue of +1+1 or 1-1, in case this Clifford is indeed stabilized by that particular Pauli operator, or to 0. For each Pauli operator, there are exponentially more Clifford states whose expectation value is 0 compared to those whose expectation value is either +1+1 or 1-1. Many common VQAs use Pauli expectation values to evaluate their cost function, as it can be implemented very efficiently and easily on a quantum computer.
The main goal of this research is to define novel methods to improve VQA ansatzes and initial states, leading to better accuracy and faster convergence rates. This goal is achieved by two novel methods:
  • Describing a new method to sample and explore VQAs’ problem space, using the DQES method we define later. This method allows one to understand the problem’s Hilbert space better, creating better guesses for initial parameter choices, and designing methods to avoid barren plateaus and local minima. This method is described in Section II.2. Examples of using this method are presented in Section III.
  • Describing an alternative heuristic method to explore the VQA’s problem space, using an efficient approach. This method aims to bring a heuristic equivalent of the above method for larger problem instances while remaining computationally feasible. This method is presented in Section II.4. Examples of using this method are also presented in Section III.
I.1 Mutually Unbiased Bases
Mutually unbiased bases (MUBs) are orthonormal bases in a Hilbert space that provide complementary and non-redundant information about a quantum state if measured [50, 56, 6, 18]. When a pure state is prepared in one basis, a measurement in any other basis of the MUBs provide no information about the prepared state.
@@ I.1 Mutually Unbiased Bases
II Methods
VQAs search for the optimal solution of the given problem using the search space imposed by the chosen ansatz. The Hilbert space of the problem is determined only by the problem definition and encoding and is not subject to changes by replacing the ansatz. The dimension of this Hilbert space is exponential in the number of qubits required to encode the problem into a quantum computer. Finding the optimal solution, or even a good approximation to it in such a huge space is usually a very hard task. In many search spaces, the gradients of the cost function become exponentially small in most areas of the search space, a phenomenon called barren plateaus, and most of the states that have significant gradients are concentrated in a small area [35, 3]. The concentration of most of the cost function values that have a significant difference between them in a small area makes the search hard even for gradient-free optimizers [2, 39].
@@ II Methods
In our research, we propose methods to learn and gain insights about the values and structure of the cost function in the problem’s Hilbert space, in a way that is agnostic to the ansatz that will be used to try to solve the problem. Using such methods can help researchers design ansatzes, optimization algorithms, and initial guesses that can avoid barren plateaus.
@@
II.1 Discretized Quantum Spanning Search
@@
Here we define discretized quantum spanning search (DQSS) which is useful for small size problems.
The cost function is usually defined using parameterized states $|{\psi(\vec{\theta})}\rangle$, produced by the ansatz. In the general case, such a search can be ansatz agnostic. When disregarding the ansatz, we replace the parameterized states with a chosen set of states from the cost function’s domain.
Definition 3.
Discretized Quantum Spanning Search (DQSS). Choose a set of states $\left\{|{\psi_{1}}\rangle\,|{\psi_{2}}\rangle,\dots,|{\psi_{k}}\rangle\right\}$ that span the problem’s Hilbert space.
Evaluate the cost function of the algorithm $C(|{\psi_{i}}\rangle)$ for $i\in\{1\ldots k\}$.
Note that in order to span the space, $k$ must be greater or equal to the dimension of the problem’s Hilbert space. For values of $k$ larger than the dimension of the problem’s Hilbert space, the spanning is considered over-complete.
@@
NP problems can be solved in polynomial time by a classical computer once a relevant and correct hint is given from an external source. In a similar way, the QMA complexity class contains problems that can be solved by a quantum computer in polynomial time once a relevant hint is given by an external source.
As explained in appendix C, in VQE, a clever scientist chooses an ansatz and guesses cleverly or randomly an initial state, from which the evaluation and optimization start. We notice that VQE resembles QMA, only instead of using the correct hint (as in QMA), a process of guessing and optimizing is used. When a classical or quantum computer tries to solve an NP problem of a small size, it is possible to test all possible classical hints via exhaustive search, gain knowledge on the structure of the problem, and design heurstics for larger size problems.
@@
As one cannot try all states of the continuous quantum space as possible hints, we look for an approximated discretized method that can resemble the classical exhaustive search.
@@
Searching an exponential number of different states, sampled from the problem’s relevant Hilbert space, for small size problems, may help design heuristics that can then be used for larger-size problems. Specifically, it may help in studying the energy landscape and energy curves of problems, hence the locations of barren plateaus and local minima.
II.2 MUBs for Exploring The Cost Function Landscape
@@
The states of all of the bases in a full set of MUBs, which we call “MUB states”, have unique properties that make them good for extracting information for purposes such as state tomography, and process tomography. In a similar manner, we suggest that these states can be used to extract information about the landscape of a cost function of a given problem, if used appropriately. A full set of MUB states can be a good candidate as states for a DQSS for small-size problems. They are an exponential number of states that are spread all over the Hilbert space, and the information learned from them is, in some sense, mutually exclusive.
@@
By measuring the cost function value of a given problem at all these states, one can get evenly-spread (throughout the Hilbert space) information about the structure of the cost function. Such information can be properties of states that give similar values (hinting that there could be barren plateaus areas in the cost function landscape), or properties of states that have a larger variety of results (pointing to small areas that might be interesting to search for the optimal result). It is important to note that such a search is done directly on the cost function landscape, therefore it can be agnostic to the ansatz used.
We define a specific method to do DQSS, by relying on a specific overcomplete spanning that uses all MUB-states of a chosen full MUB set. We call our specific method Discretized Quantum Exhaustive Search (DQES).
Definition 4.
Discretized Quantum Exhaustive Search (DQES). For a problem encoded with $n$ $p$-dimensional qudits with $p$ being a prime number, choose a full MUB set over the $p^{n}$-dimensional Hilbert space $\bm{B}=\{B_{1},\ldots,B_{P^{n}+1}\}$.
Evaluate the cost function of the algorithm $C(|{\psi^{j}_{i}}\rangle)$ for each state $i$ in each base $B_{j}$ which is part of the chosen MUB set, $i\in\{1\ldots P^{n}\},j\in\{1\ldots P^{n}+1\}$.
For the case of $n$ qubits, so that there are $(2^{n}+1)\times 2^{n}$ MUB states, we present an example of DQES in Algorithm II.2, where $j$ runs from zero to $2^{n}+1$ and $i$ runs from zero to $2^{n}$.
{algorithm}
Discrete Quantum Exhaustive Search (DQES) for $n$ qubits
1:Input: A cost function $C(|{\psi}\rangle)$ for a problem that can be defined over a $2^{n}$-dimensional Hilbert space, and a full MUB set $\bm{B}=\{B_{1},\ldots,B_{2^{n}+1}\}$
2:Define an $n$-qubit register $R$.
3:for all $B_{j}\in\bm{B}$ do
4:  for all $|{\psi^{j}_{i}}\rangle\in B_{j}$ do
5:   Set $R$’s state to be $|{\Psi}\rangle=|{\psi^{j}_{i}}\rangle$.
6:   Calculate $E_{j,i}=C(|{\Psi}\rangle)$.
7:   Save the result $(j,i,E_{j,i})$.
8:  end for
9:end for
10:Output: the collection of all $(j,i,E_{j,i})$ tuples.
In this research we stick to cases in which three MUBs are fully built from product states (it is possible for $n$ qubits, for any $n\geq 1$ [30]).
As the task of evaluating the ground state energy of a Hamiltonian (named the $k$-local Hamiltonian Problem) is Quantum Merlin Arthur (QMA)-complete [10, 27] and finding the ground state energy of Ising model Hamiltonians is NP-hard [34], our method enables obtaining intuition on NP-hard problems as well as on QMA-complete and QMA-hard problems.
We present an example of the procedure of the DQES method for learning the energy landscape of a VQE problem in Algorithm II.2. In this example, the cost function is the expectation value of the ansatz state over the Hamiltonian defining the problem. The results of the method provide an energy landscape of the problem. One then chooses all or some of the output tuples as the initial point of a variational optimization process of the VQE, in order to obtain energy curves.
{algorithm}
Discrete Quantum Exhaustive Search (DQES) for VQE
1:Input: a Hamiltonian $\mathbb{H}$ defined over an $n$-qubit register, and a full MUB set $\bm{B}=\{B_{1},\ldots,B_{2^{n}+1}\}$
2:Define an $n$-qubit register $R$.
3:for all $B_{j}\in\bm{B}$ do
4:  for all $|{\psi^{j}_{k}}\rangle\in B_{j}$ do
5:   Set $R$’s state to be $|{\Psi}\rangle=|{\psi^{j}_{k}}\rangle$.
6:   Calculate $E_{j,k}={\langle{\Psi}|}H|{\Psi}\rangle$.
7:   Save the result $(j,k,E_{j,k})$.
8:  end for
9:end for
10:Output: the collection of all $(j,k,E_{j,k})$ tuples.
II.3 Circuit Representation of MUB States
@@
The computation space of a quantum device consisting of $n$ qubits is of dimension $2^{n}$. It has been shown [6, 56, 29, 18] that for such a space there exists a set of $2^{n}+1$ orthonormal bases which are mutually unbiased. Finding such a set of mutually unbiased bases for any number of qubits is not a trivial task. Few methods have been proposed to find a full set of unbiased bases, containing $2^{n}+1$ orthonormal bases [30, 8, 6, 56]. When using those methods, the computational cost of finding orthonormal bases which are part of a full MUB set is exponential in the number of qubits. For small values of $n$, we noticed one research that found efficient ways to implement these MUBs in the lab using adjusted measurement equipment [23]. In order to perform DQES, one needs to have a quantum circuit that prepares each state in those bases. Known algorithms to find a quantum circuit that creates an arbitrary given quantum state take exponential time to compute in the worst case and might produce a very deep quantum circuit, therefore usually cannot be used in the context of VQA.
@@
Here we investigate two different circuit representations of MUB states, both sufficiently shallow for our demonstrations. In the first representation, each basis in the MUB set is defined by a quantum circuit that transforms a state in the computational basis to a state in that basis. We have manually found such shallow circuits for each basis in a full set of MUBs of 2 and 3 qubits, see Fig. 1.
This method provides only a discrete map, yet it may be used for finding an optimal initial state (or a set of states), that can then be used as the initial state(s) of a parameterized ansatz controlled by an optimizer as in all usages of VQA.
@@
In contrast with the former method, the second representation relies directly on a chosen ansatz circuit. In this representation, each MUB state is defined by the parameter vector of an ansatz which prepares the state, hence each MUB state can then directly be used by utilizing the optimizer to modify the parameters. When using this method, the chosen ansatz must be expressive enough to allow the preparation of each of the states of the MUBs. In Fig. 1 (c), we give a hardware-efficient ansatz circuit for generating 2-qubit MUBs. When acting on the computational state $|{i}\rangle$, it gives the i’th element of the MUB basis. Note that the circuit in 1 (c) can be used to generate all MUB states directly by adding a single layer of NOT gates (where needed) at the start of the circuit, if the initial state is always the all-zeros state.
@@
II.4 Efficient Partial Discretized Quantum Exhaustive Search
Another tool we suggest and demonstrate via simple examples is an efficient partial DQES. Such a search may also provide heuristics and potentially lead to exact algorithms for solving larger-size problems, by finding easy subcases or better than classical approximations.
The DQES method searches an exponential (in problem size) number of states. In this subsection we suggest an alternative method that relies on a polynomial number of states. To obtain a scalable method we let some qubits be much more dominant than others in the cost function of a VQA problem, yet we do not assume that the dominance of each qubit is known a priori. We suggest a method to find possible initial states by using a specific set of partial-MUB states.
A set of Partial-MUB states is defined as taking a MUB set over $K\ll n$ qubits, encoding the state of the MUB set into all possible allocations of $K$ indexes in an $n$-qubit register, and keeping all other qubits in some trivial state (WLOG $|{0}\rangle$).
For small values of $K$ we may generate all such MUB states. For the scalability proof we shall not assume that $K$ is small.
For a constant $K$, The number of different options to choose the qubits splitting is ${{n}\choose{K}}=O\left(n^{K}\right)$. The number of MUB states for $K$ qubits is $(2^{K}+1)\cdot 2^{K}=(2^{K})^{2}+2^{K}$, leaving us with a total number of
${{n}\choose{K}}\cdot((2^{K})^{2}+2^{K})=O(n^{K}\cdot 4^{K})=O(n^{K})$ (2)